The market price of a 11.00-year STRIPS is $347.00 The yield to maturity is ____%.
Answer #1

Hi

Yield to maturity of STRIPS is equivalent to yield to similar zero coupon bonds.

So Price of zero coupon bond = Par Value/(1+YTM)^n

Here Price = $347

Par Value = $1000

time n = 11*2 = 22 (since bonds are paid semi annually)

so

347 = 1000/(1+YTM)^22

1+YTM = (1000/347)^(1/22)

1+YTM = 1.049286

YTM = 4.9286%

So annual YTM = 2*4.9286 = 9.86%

Thanks